OpenStudy (anonymous):

Which will result in a perfect square trinomial? (3x – 5)(3x – 5) (3x – 5)(5 – 3x) (3x – 5)(3x + 5) (3x – 5)(–3x – 5) Thank youu in advance c:

OpenStudy (anonymous):

@dayakar c:

OpenStudy (anonymous):

@mathmath333 c:

OpenStudy (dayakar):

\[(3x-5)(3x-5)=\left( 3x-5 \right) ^{2}\]perfect square

OpenStudy (dayakar):

if u expand it ,u get trinomial,can u do that

OpenStudy (dayakar):

r u there

OpenStudy (anonymous):

what do you mean?

OpenStudy (dayakar):

\[\left( a-b \right)^{2}=a ^{2}-2ab+b^{2}\]

OpenStudy (dayakar):

do u understand above identity

OpenStudy (dayakar):

\[\left( 3x-5 \right)^{2}=\left( 3x \right) ^{2}-2*3x*5+5^{2}\]

OpenStudy (dayakar):

\[=9x^{2}-30x+25\]

OpenStudy (anonymous):

ohhh yes i do :) so when its aperfect trinomal they are the same :)

OpenStudy (dayakar):

above expression contains 3 terms so, it is called a trinomial

OpenStudy (dayakar):

is it clear

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

thank for all the help :)